# Subordinators and time-space fractional diffusion equations

### v1

@@ Abstract.
Nonlocal diffusion equation, Long-time behavior, Cesàro mean, Heat kernal, subordinators.
@@ 1. Introduction
Our main focus lies in understanding the behavior of the subordinated solution $v^{E}(x,t)$, which arises from applying subordination to certain evolution partial differential equations (PDEs). Specifically, we analyze large-time dynamics by convolving the original solution with the density function $G_{t}$ associated with the subordinator $\mathbf{S}$. To extract meaningful asymptotic information, we study Cesàro mean of the subordinated solution, which offers a robust measure of its average behavior.
Understanding these subordinated dynamics is not only of mathematical interest, but is also highly relevant for modeling real-world systems. In particular, our results offer potential insight into how biological time unfolds in natural processes such as species evolution and ecological development. This perspective enables the development of more refined models in fields such as population dynamics, epidemiology, and ecosystem analysis. For more context and foundational background, we refer the reader to [19, 20].
@@
where $\mathscr{L}_{a,b}=a\Delta-b(-\Delta)^{s}$ is a mixed local-nonlocal operator, $a,b,\gamma\geq 0$, $0<\alpha\leq 1$, $0<\mathsf{s}<1$, and $\varphi(x)$ is a nonnegative initial datum. Here, $\partial_{t}^{\alpha}$ denotes the Caputo fractional derivative [8], while $(-\Delta)^{s}$ represents the fractional Laplacian, defined via the Fourier transform as
@@
Building on the framework developed in [3], we observe that in the special case $\alpha=1$, $\gamma=0$, the subordinated solution $v^{E}(x,t)$ satisfies the following fractional differential equation
$\begin{cases}\left(\mathbb{D}^{(k)}_{t}v^{E}\right)(x,t)=\mathscr{L}_{a,b}\,v^{E}(x,t),&(x,t)\in\mathbb{R}^{N}\times(0,\infty)\\ v^{E}(x,0)=\varphi(x),&x\in\mathbb{R}^{N}\end{cases}$ (1.3)
Here $\mathbb{D}^{(k)}_{t}$ denotes the differential-convolution operator defined by
@@
with kernel $\mathbf{k}\in L^{1}_{\text{loc}}(0,\infty)$ being positive and locally integrable.
@@ 2.1. Subordinators and Their Properties
These processes provide fundamental models for random time changes and jump phenomena [4, 5].
The infinite divisibility of $\mathbf{S}_{t}$ implies its Laplace transform takes the form:
@@ 2.2. Caputo Fractional Derivative
Under the stronger assumption that $u$ is absolutely continuous, the Caputo derivative admits the equivalent representation [11, Theorem 2.1, p. 92]:
@@ 2.3. Mixed Heat Semigroup
The proof appears in [21, p. 395], with the positivity result first stated without proof in [7, p. 263]. A detailed treatment can also be found in [7, Theorem 2.1].
@@ 2.4. Karamata’s Tauberian theorem
The Feller-Karamata Tauberian theorem represents a fundamental result in asymptotic analysis, providing a deep connection between the asymptotic behavior of a function and its Laplace transform. For comprehensive treatments of the general result, we refer to [6, Section 1.7] and [9, Chapter XIII, Section 5]. Below, we provide a version of this theorem that is particularly well-suited to our applications.
@@ 4. Concluding Remarks
The study of subordinated solutions to fractional diffusion equations with mixed local and nonlocal operators provides a powerful way to understand complex systems that exhibit memory effects and anomalous diffusion. By applying techniques from probability theory, asymptotic analysis, and partial differential equations (PDEs), we have analyzed the long-term behavior of these solutions, particularly through the Cesàro mean, which gives a reliable measure of their average dynamics.

### v2

# Long-Time Asymptotics for Subordinated Fractional Diffusion Equations

@@ Abstract.
Nonlocal diffusion equation, Long-time behavior, Cesàro mean, Heat kernel, subordinators.
@@ 1. Introduction
Our main focus lies in understanding the behavior of the subordinated solution $v^{E}(x,t)$, which arises from applying subordination to certain evolution partial differential equations (PDEs). Specifically, we analyze large-time dynamics by convolving the original solution with the density function $G_{t}$ associated with the subordinator $\mathbf{S}$. To extract meaningful asymptotic information, we study the Cesàro mean of the subordinated solution, which provides a robust measure of its average behavior.
Understanding these subordinated dynamics is not only of mathematical interest, but is also highly relevant for modeling real-world systems. In particular, Our results offer insight into how biological time may unfold in natural processes such as species evolution and ecological development. This perspective enables the development of more refined models in fields such as population dynamics, epidemiology, and ecosystem analysis. For more context and foundational background, we refer the reader to [20, 21].
@@
where $\mathscr{L}_{a,b}=a\Delta-b(-\Delta)^{s}$ is a mixed local-nonlocal operator, $a,b,\gamma\geq 0$, $0<\alpha\leq 1$, $0<\mathsf{s}<1$, and $\varphi(x)$ is a nonnegative initial datum. Here, $\partial_{t}^{\alpha}$ denotes the Caputo fractional derivative [9] while $(-\Delta)^{s}$ represents the fractional Laplacian, defined via the Fourier transform as
@@
In the particular case when $\alpha=1$ and $\gamma=0$, it is known—under suitable assumptions (see [14, 15])—that the subordinated solution $v^{E}(x,t)$ satisfies the following fractional evolution equation:
$\begin{cases}\left(\mathbb{D}^{(k)}_{t}v^{E}\right)(x,t)=\mathscr{L}_{a,b}\,v^{E}(x,t),&(x,t)\in\mathbb{R}^{N}\times(0,\infty),\\[4.0pt] v^{E}(x,0)=\varphi(x),&x\in\mathbb{R}^{N}.\end{cases}$ (1.3)
Here, $\mathbb{D}^{(k)}_{t}$ denotes a convolution-type derivative defined by
@@
where the kernel $\mathbf{k}\in L^{1}_{\mathrm{loc}}(0,\infty)$ is assumed to be positive and locally integrable.
@@ 2.1. Subordinators and Their Properties
These processes provide fundamental models for random time changes and jump phenomena [4, 6, 5].
The infinite divisibility of $\mathbf{S}_{t}$ implies that its Laplace transform takes the form:
@@ 2.2. Caputo Fractional Derivative
Under the stronger assumption that $u$ is absolutely continuous, the Caputo derivative admits the equivalent representation [12, Theorem 2.1, p. 92]:
@@ 2.3. Mixed Heat Semigroup
The proof appears in [22, p. 395], with the positivity result first stated without proof in [8, p. 263]. A detailed treatment can also be found in [8, Theorem 2.1].
@@ 2.4. Karamata’s Tauberian theorem
The Feller-Karamata Tauberian theorem represents a fundamental result in asymptotic analysis, providing a deep connection between the asymptotic behavior of a function and its Laplace transform. For comprehensive treatments of the general result, we refer to [7, Section 1.7] and [10, Chapter XIII, Section 5]. Below, we provide a version of this theorem that is particularly well-suited to our applications.
@@ 4. Concluding Remarks
The study of subordinated solutions to fractional diffusion equations with mixed local and nonlocal operators provides a useful framework for understanding complex systems that exhibit memory effects and anomalous diffusion. By applying techniques from probability theory, asymptotic analysis, and partial differential equations (PDEs), we have analyzed the long-term behavior of these solutions, particularly through the Cesàro mean, which gives a reliable measure of their average dynamics.